OpenStudy (anonymous):

What is the value of the expression 2^4 – 3^3? A. 7 B. 1 C. –11 D. –1

jimthompson5910 (jim_thompson5910):

What is the value of 2^4 ?

OpenStudy (anonymous):

Remember 2^4 = 2*2*2*2

OpenStudy (anonymous):

2^4=2*2*2*2 3^3=3*3*3 16-27=

OpenStudy (anonymous):

C.-11

OpenStudy (anonymous):

Give op the chance to solve it.

OpenStudy (anonymous):

its -11 i did the math my self i didnt give up @iambatman

OpenStudy (anonymous):

Good job @needhelpsuperfast !

OpenStudy (anonymous):

thanks!